# HERACLES++: a multi-dimensional Eulerian code for exascale computing

### v1

@@ 5.5.1 Setup
The 3-D simulation with HERACLES++ was initialized with the same 1-D input described in Section 5.2. This 1-D profile was remapped onto the same radial grid and copied across all angular directions $(\theta,\phi)$ both in the range $[\pi/4,3\pi/4]$. We used the same setup as in Section 5.2 and add reflexive boundary conditions for the angular directions. We chose a grid with $[n_{r},n_{\theta},n_{\phi}]=[1024,512,512]$. Simulations using other resolutions in both radius and angle are presented in Appendix A. In the preliminary tests for the 3-D simulations, we employed a variety of initial perturbations of the density and velocity to act as seeds for the Rayleigh-Taylor instability. We find, however, that the instability grows from numerical noise and also aided by the small-scale high-frequency oscillations present in the V1D input (i.e., in the post-shock region at the highest velocities). So, no initial perturbation of the fluid variables was introduced for the tests presented in this work.
Figure 12: Evolution of the angle-averaged density (top), velocity dispersion (middle) and composition (bottom; the solid [dashed] line corresponds to H [He]) from the initial time until shock emergence and versus Lagrangian mass. We show the same epochs as in Fig. 11. In the top panel, the final density obtained in the 1-D counterpart is shown as a dashed line. (See Section 5.5.2 for discussion.)
Figure 13: Evolution of the angular properties of the quantity $\log(\rho/\langle\rho\rangle$) at a fixed Lagrangian mass (4.5 $M_{\odot}$). In the top row and from left to right, we show colormaps of radial sections through the 3-D domain at 127, 323, 500 and 635 $R_{\odot}$ and at times 1, 4, 8 and 12 ×$\times 10^{4}$ s, respectively. In the bottom row, we show the fractional shell volume per normalized density bin for each of the shells shown in the top row.
Figure 14: 3-D rendering of the helium mass fraction in the supernova test for two different viewing angles. The time is $12\times 10^{4}$ s.
@@ 5.5.2 Results
In this section, we present a 3-D simulation of the perturbed 1-D explosion model presented in Section 5.2. To facilitate the discussion of the results, we first present the evolution of various fluid quantities within meridional slices through the 3-D domain. Then, we also discuss angle-averaged quantities for the same epochs but shown versus Lagrangian mass rather than radius. In the following, “<>absent<>< >” is meant as an average over the polar and azimuthal angles at a given radius or Lagrangian mass.
Figure 11 shows the radial velocity, the logarithm of the normalized density, the polar velocity and the He mass fraction in the $\phi=\pi/2$ plane at 1, 4, 8, and $12\times 10^{4}$ s – each quadrant corresponds to a time ordered chronologically from twelve noon. These slices are representative of the overall evolution of the 3D ejecta and other choices for these slices (i.e., a different azimuth, or cuts instead in the $(r,\phi)$ plane exhibit similar properties).
As in 1-D, we see that a pair of reverse and forward shocks form soon after the start of the simulation. However, in contrast to the 1-D simulation, fluctuations of the velocity, the density, as well as the composition, quickly appear throughout the region in between the shocks. These shocks are clearly visible in Fig. 11a, bracketing a turbulent flow which is located between about 450 and 900 $R_{\odot}$ at $12\times 10^{4}$ s. The reverse shock remains roughly spherically symmetric but the forward shock is corrugated, showing variations in extent at the 5–10 % level. The structured, unstable flow extends all the way to the forward shock. This probably arises from the strong deceleration of the forward shock as it propagates through the extended red-supergiant star envelope (compare the first and last quadrants in Fig. 11a; see also Fig. 8).
Figure 11b indicates that the density variations are maximum at the forward shock, reaching a factor of a few the angular average at that radius, and to a lesser extent at the reverse shock. This is partially an artifact from showing the density normalized to the average and the fact that the forward shock is corrugated. Within the shocked region, the density enhancements appear as filaments surrounded by more extended, lower density regions that are close to the average density.
@@
The morphology of the instability-generated structures is more vividly illustrated by the He distribution, which is the dominant element in the He-rich shell (at the 95 % level) and the second-most abundant in the H-rich shell (at the 28 % level). In Fig. 11d, we see the development of He-rich fingers, plumes, and mushrooms into the H-rich envelope, stretching out as time passes, and reaching out to the forward shock at the last time shown. These structures are well resolved in our calculations, which use 512 polar and azimuthal zones for a full coverage in each direction of ninety degrees(see Appendix A for a comparison of the structure with different radial and angular resolutions.). The He mass fraction in the most radially-elongated structures is, however, much smaller than the original He mass fraction in the He-rich shell. These structures have probably expanded laterally and mixed with the H-rich material, which may in part arise from numerical diffusion.
@@
To characterize more quantitatively the lateral structure in the 3-D simulation, Fig. 13 shows the properties of a slice of the quantity $\log(\rho/\langle\rho\rangle)$ at different radii and times, and starting at 104 s. We select a slice at 120 $R_{\odot}$ initially (around a Lagrangian mass of 4 $M_{\odot}$), which is located just below the forward shock at 137 $R_{\odot}$ (the choice of radius/time is made such that we track the same slice of the ejecta as it advects outwards). The top row shows colormaps of the density variations relative to the polar and azimuthal angles, whereas the bottom row shows histograms of the fractional volume (within that slice or shell) per bin in $\log(\rho/\langle\rho\rangle)$. At 104 s, structures are within 50 % of the average , with naturally a larger volume occupied by the underdense material (the histogram is slightly skewed relative to the mean density). At the next epoch, the density contrast between underdense and overdense material has grown, with the density maximum reaching four times the average density. The histogram is now highly skewed with about 65 % of the volume occupied by material about 50 % below the average density. The structures are randomly distributed in angle (and remain so at all times). As time progresses and the ejecta expand further, the histogram becomes more symmetric and the entire material is found within 50 % of the mean density. There is thus only a modest residual density contrast in the shocked envelope by the time the shock emerges from the progenitor star.
Finally, we show a 3-D rendering of the He mass fraction in Fig. 14 for two lines of sight looking radially inwards from outside as well as from the side. This gives an alternative view of the random distribution of the advected He-rich structures into the H-rich envelope of the progenitor star as well as the filamentary nature of the most He-rich material advecting out. The low He-mass fraction of the structures located at the largest radii suggests some level of numerical diffusion.
@@ 6 Conclusion
In this paper, we have presented the new multidimensional hydrodynamics code HERACLES++, which is designed for computing at exascale, with a parallelization that works both for CPUs and GPUs, and equipped with a good portability across a variety of architectures. The code is written in C++ and makes use of the Kokkos library. HERACLES++ is a Eulerian code that works in Cartesian and spherical coordinates and in 1-D, 2-D, or 3-D. Passive scalars are employed to handle the evolution of a nonuniform initial composition. The mesh may be irregular, adopting for example a combination of a linear grid and a logarithmic grid. The code currently handles constant, point-mass, and interior-mass gravity. Two equations of state are possible with either a pure ideal gas or a mixture of ideal gas and radiation. The code accepts a user-supplied heating or cooling term in order to treat astrophysical conditions such as radioactive decay in supernova ejecta or radiative losses from the photosphere of a star. The code was benchmarked with a variety of tests. We show here the results for some usual 1-D tests (e.g., a 1-D Sod shock tube or a 2-D implosion) as well as a 2-D Cartesian Kelvin-Helmholtz test, a 3-D Cartesian Rayleigh-Taylor test, and a 3-D Spherical off-center blast wave test. Performance tests (see Appendix B) were also carried out and reveal a satisfactory scalability of the code.
@@
We also conducted a more astrophysics-oriented test with the study of the propagation of a supernova shock through the envelope of a red-supergiant star, using as initial conditions a model of an exploding massive star (Ercolino et al., 2024; Dessart et al., 2024). This is a useful test, which has been studied in detail in the context of SN 1987A by Fryxell et al. (1991) and Mueller et al. (1991). In 1-D, HERACLES++ produces results for the evolution of the density, temperature, or velocity that are in excellent agreement with those obtained with the 1-D Lagrangian (radiation) hydrodynamics code V1D (Livne, 1993). Because this configuration is Rayleigh-Taylor unstable, we repeat this calculation in 3-D and evolve the model until the supernova shock emerges at the surface of the progenitor star. We find that the instability develops from the earliest times, eventually extending throughout the shocked region, and leading to filamentary structures or fingers, corresponding to variations in density of a factor of a few at several 104 s and dropping to a few tens of percent by the time of shock emergence. The instability also erases the density jump that persists in 1-D counterparts at the base of the H-rich layers (i.e., the interface between the H-rich envelope and the He core in the pre-supernova star). There is dispersion in velocity both in the radial and angular directions, which may survive as turbulence in the ejecta after shock emergence. The most vivid illustration of this instability may be seen in the impact it has on the composition, with advection of H-rich gas down to the H-free innermost regions and advection of He-rich gas from the He-rich shell throughout the H-rich envelope. This behavior reproduces qualitatively previous studies on similar shocked envelopes.
@@
This benchmarking test will serve as a basis of a forthcoming study on the shock propagation in supergiant stars differing in H-rich envelope mass as well as the long-term influence of ${}^{56}\rm{N}$i-decay heating in such shocked envelopes. Future work will also be to extend HERACLES++ to a full radiation-hydrodynamics code.
@@ Appendix A Resolution study
Figure 15: Impact of the radial resolution on the results of the 3-D supernova shock simulation presented in Section 5.5.2. We show colormaps of the quantity $\log(\rho\langle\rho\rangle)$ in a meridional slice at $\phi=\pi/2$ and corresponding to a time of $10^{5}$ s. Clockwise from twelve noon, we show the results for increasing radial resolution by varying $n_{r}$ from 256 to 512, 768, and 1024 – the angular resolution is kept fixed with $n_{\theta}=n_{\phi}=256$.
@@
Figure 18: Same as Fig. 16 but now for the simulations differing in angular resolution and shown in Fig. 17.
We now present the results from resolution tests. We use the same setup as in Section 5.5 but now vary the number of radial zones or the number of angular zones to assess the impact on the R-T instability-generated structures at the time near shock emergence.
@@
Fig. 17 is a counterpart of Fig. 15 but now showing the results for both $n_{\theta}$ and $n_{\phi}$ increased from 128 to 256, 512, and 768, while the number of radial zones is kept fixed at 768. In contrast to the radial-resolution dependence, the models with higher angular resolution produce a pair of reverse and forward shocks that reach smaller and bigger radii, respectively (see also Fig. 18. The Rayleigh-Taylor instability-generated structures have a smaller angular scale at higher angular resolution although they are clearly present and to some extent resolved even with the lower resolution employed here. This low resolution corresponds to 0.7 deg and is already greater than typically employed in 3-D simulations of core-collapse supernovae (for example, an angular resolution of 1 deg is used in Wongwathanarat et al. 2015).
This study helps us define the size of the simulation for the astrophysical test of the code (Section 5). We needed to balance good resolution of the physics while performing simulations of reasonable size, given the available computational time, resources, and our ability to analyze them. We chose to work with a grid having $n_{r}=1024$ and $n_{\theta}=n_{\phi}=512$, corresponding to a total of about $2.7\times 10^{8}$ cells.
@@ Appendix B Performance evaluation
For this analysis, we used the 3-D Rayleigh-Taylor test in Cartesian coordinates discussed in Section 4.2). Initially, we do not employ any passive scalar. We maximize the size of the problem with one GCD, which is $5\times 10^{7}$ cells. After determining the maximum size on a single GPU, we repeated the calculation, each time doubling both the size of the problem and the number of GPUs. The results are shown in Fig. 19. The ideal weak scaling is calculated based on the performance of a single MPI process and is represented as a line from this assumption. Our results slightly differ from this line.
We then compared with a test based on the configuration for the propagation of a supernova shock in a stellar envelope and discussed in Section 5. The performance is now significantly lower with $1.3\times 10^{10}$ cells s-1 compared to $2.3\times 10^{10}$ cells s-1 in the Cartesian case and 512 MPI processes. This decrease is due to the additional geometrical terms appearing in the equations when spherical coordinates are used as well as the use of an EoS with a mixture of ideal gas and radiation. The use of passive scalars also comes with a sizable computational cost, further reducing the performance in HERACLES++.
We have also tested HERACLES++  on a variety of other architectures including Nvidia GPU’s from the IDRIS computing center and various kind of CPUs (from standard notebook to supercomputers). Thanks to the usage of Kokkos, HERACLES++  can run efficiently on all these systems just by changing the compiling options.
@@ Appendix C Additional hydrodynamic tests
Figure 20: Results for the 1-D Sod shock tube test. We show the profiles at $t=0.2$ for the density (top), the velocity (middle), and the pressure (bottom) obtained with HERACLES++ (solid) and predicted analytically (dashed).
Figure 21: Evolution of the passive scalar $f_{\rm x}$ for the 2-D Rayleigh-Taylor instability test. From left to right, we show the results at $t=0$, 1.5, and 3. (see Section C.2 for details.)

### v2

@@ 5.5.1 Setup
The 3D simulation with HERACLES++ was initialized with the same 1D input described in Sect. 5.2. This 1D profile was remapped onto the same radial grid and copied across all angular directions $(\theta,\phi)$ both in the range $[\pi/4,3\pi/4]$. We used the same setup as in Sect. 5.2 and added reflexive boundary conditions for the angular directions. We chose a grid with $[n_{r},n_{\theta},n_{\phi}]=[1024,512,512]$. Simulations using other resolutions in both radius and angle are presented in Appendix A. In the preliminary tests for the 3D simulations, we employed a variety of initial perturbations of the density and velocity to act as seeds for the Rayleigh-Taylor instability. We find, however, that the instability grows from numerical noise and is also aided by the small-scale high-frequency oscillations present in the V1D input (i.e., in the post-shock region at the highest velocities). So, no initial perturbation of the fluid variables was introduced for the tests presented in this work.
Figure 12: Evolution of the angle-averaged density (top), velocity dispersion (middle), and composition (bottom; the solid [dashed] line corresponds to H [He]) from the initial time until shock emergence and versus Lagrangian mass. We show the same epochs as in Fig. 11. In the top panel, the final density obtained in the 1D counterpart is shown as a dashed line. (See Sect. 5.5.2 for discussion.)
Figure 13: Evolution of the angular properties of the quantity $\log(\rho/\langle\rho\rangle$) at a fixed Lagrangian mass (4.5 $M_{\odot}$). In the top row and from left to right, we show color maps of radial sections through the 3D domain at 127, 323, 500, and 635 $R_{\odot}$ and at times of 1, 4, 8, and 12 ×$\times 10^{4}$ s, respectively. In the bottom row, we show the fractional shell volume per normalized density bin for each of the shells shown in the top row.
Figure 14: 3D rendering of the helium mass fraction in the SN test for two different viewing angles. The time is $12\times 10^{4}$ s.
@@ 5.5.2 Results
In this section, we present a 3D simulation of the perturbed 1D explosion model presented in Sect. 5.2. To facilitate the discussion of the results, we first present the evolution of various fluid quantities within meridional slices through the 3D domain. Then, we also discuss angle-averaged quantities for the same epochs but shown versus the Lagrangian mass rather than the radius. In the following, “<>absent<>< >” is meant as an average over the polar and azimuthal angles at a given radius or Lagrangian mass.
Figure 11 shows the radial velocity, the logarithm of the normalized density, the polar velocity, and the He mass fraction in the $\phi=\pi/2$ plane at 1, 4, 8, and $12\times 10^{4}$ s – each quadrant corresponds to a time ordered chronologically from twelve noon. These slices are representative of the overall evolution of the 3D ejecta and other choices for these slices (i.e., a different azimuth, or cuts instead in the $(r,\phi)$ plane exhibit similar properties).
As in 1D, we see that a pair of reverse and forward shocks form soon after the start of the simulation. However, in contrast to the 1D simulation, fluctuations in the velocity, the density, and the composition quickly appear throughout the region in between the shocks. These shocks are clearly visible in Fig. 11a, bracketing a turbulent flow that is located between about 450 and 900 $R_{\odot}$ at $12\times 10^{4}$ s. The reverse shock remains roughly spherically symmetric but the forward shock is corrugated, showing variations in extent at the 5–10 % level. The structured, unstable flow extends all the way to the forward shock. This probably arises from the strong deceleration of the forward shock as it propagates through the extended RSG star envelope (compare the first and last quadrants in Fig. 11a; see also Fig. 8).
Figure 11b indicates that the density variations are maximum at the forward shock, reaching a factor of a few the angular average at that radius, and to a lesser extent at the reverse shock. This is partially an artifact from showing the density normalized to the average and the fact that the forward shock is corrugated. Within the shocked region, the density enhancements appear as filaments surrounded by more extended, lower-density regions that are close to the average density.
@@
The morphology of the instability-generated structures is more vividly illustrated by the He distribution, which is the dominant element in the He-rich shell (at the 95 % level) and the second-most abundant in the H-rich shell (at the 28 % level). In Fig. 11d, we see the development of He-rich fingers, plumes, and mushrooms into the H-rich envelope, stretching out as time passes, and reaching out to the forward shock at the last time shown. These structures are well resolved in our calculations, which use 512 polar and azimuthal zones for full coverage in each direction of ninety degrees (see Appendix A for a comparison of the structure with different radial and angular resolutions). The He mass fraction in the most radially elongated structures is, however, much smaller than the original He mass fraction in the He-rich shell. These structures have probably expanded laterally and mixed with the H-rich material, which may in part arise from numerical diffusion.
@@
To characterize more quantitatively the lateral structure in the 3D simulation, Fig. 13 shows the properties of a slice of the quantity $\log(\rho/\langle\rho\rangle)$ at different radii and times, and starting at 104 s. We selected a slice at 120 $R_{\odot}$ initially (around a Lagrangian mass of 4 $M_{\odot}$), which is located just below the forward shock at 137 $R_{\odot}$ (the choice of the radius at the different times is made such that we track the same slice of the ejecta as it advects outward). The top row shows color maps of the density variations relative to the polar and azimuthal angles, whereas the bottom row shows histograms of the fractional volume (within that slice or shell) per bin in $\log(\rho/\langle\rho\rangle)$. At 104 s, structures are within 50 % of the average, with naturally a larger volume occupied by the underdense material (the histogram is slightly skewed relative to the mean density). At the next epoch, the density contrast between underdense and overdense material has grown, with the density maximum reaching four times the average density. The histogram is now highly skewed with about 65 % of the volume occupied by material about 50 % below the average density. The structures are randomly distributed in angle (and remain so at all times). As time progresses and the ejecta expand further, the histogram becomes more symmetric and the entire material is found within 50 % of the mean density. There is thus only a modest residual density contrast in the shocked envelope by the time the shock emerges from the progenitor star.
Finally, we show a 3D rendering of the He mass fraction in Fig. 14 for two lines of sight looking radially inward from outside as well as from the side. This gives an alternative view of the random distribution of the advected He-rich structures into the H-rich envelope of the progenitor star as well as the filamentary nature of the most He-rich material advecting out. The low He-mass fraction of the structures located at the largest radii suggests some level of numerical diffusion.
@@ 6 Conclusion
In this paper, we have presented the new multidimensional hydrodynamics code HERACLES++, which is designed for computing at exascale, with a parallelization that works both for CPUs and GPUs, and equipped with a good portability across a variety of architectures. The code is written in C++ and makes use of the Kokkos library. HERACLES++ is a Eulerian code that works in Cartesian and spherical coordinates and in 1D, 2D, or 3D. Passive scalars are employed to handle the evolution of a nonuniform initial composition. The mesh may be irregular, adopting for example a combination of a linear grid and a logarithmic grid. The code currently handles constant, point-mass, and interior-mass gravity. Two EoSs are possible with either a pure ideal gas or a mixture of ideal gas and radiation. The code accepts a user-supplied heating or cooling term in order to treat astrophysical conditions such as radioactive decay in SN ejecta or radiative losses from the photosphere of a star. The code was benchmarked with a variety of tests. We show here the results for some usual 1D tests (e.g., a 1D Sod shock tube or a 2D implosion) as well as a 2D Cartesian Kelvin-Helmholtz test, a 3D Cartesian Rayleigh-Taylor test, and a 3D Spherical off-center blast wave test. Performance tests (see Appendix B) were also carried out and reveal a satisfactory scalability of the code.
@@
We also conducted a more astrophysics-oriented test with the study of the propagation of a SN shock through the envelope of a RSG star, using as initial conditions a model of an exploding massive star (Ercolino et al., 2024; Dessart et al., 2024). This is a useful test, which has been studied in detail in the context of SN 1987A by Fryxell et al. (1991) and Mueller et al. (1991). In 1D, HERACLES++ produces results for the evolution of the density, temperature, or velocity that are in excellent agreement with those obtained with the 1D Lagrangian (radiation) hydrodynamics code V1D (Livne, 1993). Because this configuration is Rayleigh-Taylor unstable, we repeated this calculation in 3D and evolved the model until the SN shock emerged at the surface of the progenitor star. We find that the instability develops from the earliest times, eventually extending throughout the shocked region, and leading to filamentary structures or fingers, corresponding to variations in density of the order of a few at several 104 s and dropping to a few tens of percent by the time of shock emergence. The instability also erases the density jump that persists in 1D counterparts at the base of the H-rich layers (i.e., the interface between the H-rich envelope and the He core in the pre-SN star). There is dispersion in velocity both in the radial and angular directions, which may survive as turbulence in the ejecta after shock emergence. The most vivid illustration of this instability can be seen in the impact it has on the composition, with advection of H-rich gas down to the H-free innermost regions and advection of He-rich gas from the He-rich shell throughout the H-rich envelope. This behavior qualitatively reproduces previous studies on similar shocked envelopes.
@@
This benchmarking test will serve as a basis of a forthcoming study on the shock propagation in supergiant stars differing in H-rich envelope mass as well as the long-term influence of ${}^{56}\rm{N}$i-decay heating in such shocked envelopes. Future work will also extend HERACLES++ to a full radiation-hydrodynamics code.
@@ Appendix A Resolution study
Figure 15: Impact of the radial resolution on the results of the 3D SN shock simulation presented in Sect. 5.5.2. We show color maps of the quantity $\log(\rho/\langle\rho\rangle)$ in a meridional slice at $\phi=\pi/2$ and corresponding to a time of $10^{5}$ s. Clockwise from twelve noon, we show the results for increasing radial resolution by varying $n_{r}$ from 256 to 512, 768, and 1024 – the angular resolution is kept fixed with $n_{\theta}=n_{\phi}=256$.
@@
Figure 18: Same as Fig. 16 but for the simulations differing in angular resolution and shown in Fig. 17.
We now present the results from resolution tests. We use the same setup as in Sect. 5.5 but now vary the number of radial zones or the number of angular zones to assess the impact on the R-T instability-generated structures at the time near shock emergence.
@@
Figure 17 is a counterpart of Fig. 15 but now showing the results for both $n_{\theta}$ and $n_{\phi}$ increased from 128 to 256, 512, and 768, while the number of radial zones is kept fixed at 768. In contrast to the radial-resolution dependence, the models with higher angular resolution produce a pair of reverse and forward shocks that reach smaller and bigger radii, respectively (see also Fig. 18. The Rayleigh-Taylor instability-generated structures have a smaller angular scale at higher angular resolution although they are clearly present and to some extent resolved even with the lower resolution employed here. This low resolution corresponds to 0.7 deg and is already greater than typically employed in 3D simulations of core-collapse SNe (for example, an angular resolution of 1 deg is used in Wongwathanarat et al. 2015).
This study helps us define the size of the simulation for the astrophysical test of the code (Sect. 5). We needed to balance good resolution of the physics while performing simulations of reasonable size, given the available computational time, resources, and our ability to analyze them. We chose to work with a grid having $n_{r}=1024$ and $n_{\theta}=n_{\phi}=512$, corresponding to a total of about $2.7\times 10^{8}$ cells.
@@ Appendix B Performance evaluation
For this analysis, we used the 3D Rayleigh-Taylor test in Cartesian coordinates discussed in Sect. 4.2. Initially, we do not employ any passive scalar. We maximize the size of the problem with one GCD, which is $5\times 10^{7}$ cells. After determining the maximum size on a single GPU, we repeated the calculation, each time doubling both the size of the problem and the number of GPUs. The results are shown in Fig. 19. The ideal weak scaling is calculated based on the performance of a single MPI process and is represented as a line from this assumption. Our results slightly differ from this line.
We then compared with a test based on the configuration for the propagation of a SN shock in a stellar envelope and discussed in Sect. 5. The performance is now significantly lower with $1.3\times 10^{10}$ cells s-1 compared to $2.3\times 10^{10}$ cells s-1 in the Cartesian case and 512 MPI processes. This decrease is due to the additional geometrical terms appearing in the equations when spherical coordinates are used as well as the use of an EoS with a mixture of ideal gas and radiation. The use of passive scalars also comes with a sizable computational cost, further reducing the performance in HERACLES++.
We have also tested HERACLES++  on a variety of other architectures including Nvidia GPU’s from the IDRIS computing center and various kinds of CPUs (from standard notebook to supercomputers). Thanks to the usage of Kokkos, HERACLES++  can run efficiently on all these systems just by changing the compiling options.
@@ Appendix C Additional hydrodynamic tests
Figure 20: Results for the 1D Sod shock tube test. We show the profiles at $t=0.2$ for the density (top), the velocity (middle), and the pressure (bottom) obtained with HERACLES++ (solid) and predicted analytically (dashed).
Figure 21: Evolution of the passive scalar $f_{\rm x}$ for the 2D Rayleigh-Taylor instability test. From left to right, we show the results at $t=0$, 1.5, and 3. (see Appendix C.2 for details.)